Find the rule of the quadratic function whose graph passes through the given points (one of which is the vertex). (0,10),(-4,-6),(3,43)?
I need an answer by tonight 2/29/12 please.

Answer by stanbon(75887)

Find the rule of the quadratic function whose graph passes through the given points (one of which is the vertex). (0,10),(-4,-6),(3,43)?
-----
Use each point to create an equation with variable a,b,c.
Solve the system of 3 equation to find a,b,c.
------------------------------------
ax^2 + bx + c = y
----
(0,10) gives: 0a + 0b + c = 10
----
(-4,6) gives: 16a-4b + c = -6
-----
(3,43) gives: 9a + 3b + c = 43
-------
I used the TI-84 Matrix function to get:
a = 1
b = 8
c = 10
---------------
Equation:
y = x^2+8x+10
=====================
Cheers,
Stan H.
=====================